OpenStudy (anonymous):

Simplify. Assume the divisor is not zero. (4x^2y^2 – 4xy) ÷ 4xy

OpenStudy (the_fizicx99):

4x^2y^2/4xy = 4/4 = 1 x^2/x = x y^2 / y = y 1xy or just xy

OpenStudy (the_fizicx99):

-4xy/4xy is the same as saying -1/1 it'll be -1, Anything over itself will be 1, however, because one is negative your answer will be -1 So, xy -1